A farmer has both pigs and chickens on his farm.
There are 78 feet and 27 heads. How many pigs and chickens are there?

Found 2 solutions by greenestamps, Edwin McCravy:
Answer by greenestamps(13200)



First a typical standard algebraic solution....

p = # of pigs
c = # of chickens

Each pig has four feet and each chicken has two; each pig and each chicken has one head.

The total number of heads is 27:
p+c = 27 [1]

The total number of feet is 78:
4p+2c = 78 [2]

Solve the pair of equations [1] and [2] by your favorite method. When the two equations are in this form, elimination is my preference.

Multiply equation [1] by 2 and compare to equation [2] to eliminate c and solve for p.

2p+2c = 54
4p+2c = 78
2p = 78-54 = 24
p = 24/2 = 12

The number of pigs is 12, so the number of chickens is 27-12 = 15.

ANSWER: 12 pigs, 15 chickens

And now an informal solution, done mentally and quickly.

If all 27 animals were chickens, the number of feet would be 27*2 = 54.
The actual number of feet is 78, which is 24 more than that.
Each pig has 2 more feet than each chicken, so the number of pigs is 24/2 = 12.

ANSWER: 12 pigs and 17-12 = 15 chickens

For studying math, understanding the formal algebraic solution is important.

But solving problems like this informally using logical reasoning is excellent brain exercise.


Answer by Edwin McCravy(20055)


The 27 animal heads can't be all be chicken heads, for if they were, 
they'd have only 
27x2=54
feet under them. So where did the other 
78-54=24 
feet come from?  
From the 24/2=12 pigs' right and left hind feet.  
So there were 12 pigs and 27-12=15 chickens.

Or, you can solve it this way:

The 27 animal heads can't all be pig heads for if they were, they'd have 
27x4=108 
feet under them. So how did the 108 feet get reduced by 108-78=30? 
By the 30/2=15 chicken heads' failure to have a pair of hind feet 
under them.
So there were 15 chickens and 27-15=12 pigs.  

Edwin